OpenStudy (djbreezy):

which equation in point-slope form for the given point and slope? Point:(7,4); Slope:6 A. y-7=6(x+4) B. y-4=6(x-7) C. y-4=6(x+7) D. y+4=6(x-7)

rebeccaxhawaii (rebeccaxhawaii):

well solve for y first for each equation

OpenStudy (anonymous):

First use the formula of the slope intercept form that's y =mx+b

OpenStudy (texaschic101):

y - y1 = m( x - x1) slope(m) = 6 (7,4)....x1 = 7 and y1 = 4....now just sub

OpenStudy (faiqraees):

\[\large\rm (y-y_1)= slope*(x-x_1) \text{where}(x_1,y_1) \text{ are the coordinate of any point}\]\[\large\rm \text{on the line.}\]

OpenStudy (anonymous):

The answer is B 😃

OpenStudy (jeansweets):

Y-4=6(X-7)👍

OpenStudy (djbreezy):

I dont know how to do this at all but i have 4 more questions same as this